# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 8.5. The Hacking–Keel–Yu conjecture
Putting together our results in previous sections, we finally arrive at the following theorem, proving the Hacking–Keel–Yu conjecture.
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with a maximal degeneration defined by $(P,\mathscr{P})$ as in §2.4, and let ${\mathcal{X}_{\mathscr{P}}}$ be its semistable mirror. Then, the coarse KSBA moduli space $M_{(Y,D,L)}$ admits a toric finite cover $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow\mathcal{M}_{(Y,D,L)}$, where $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the toric variety associated to the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ as in Definition 6.26.
@@
By Theorem 7.13, the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is the base of a family of KSBA stable log Calabi–Yau surfaces which by Theorem 8.10 are generically deformation equivalent to $(Y,D,L)$. Hence, by the universal property of the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ we obtain a map of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$, and so an induced map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to M_{(Y,D,L)}$ at the level of coarse moduli spaces.
To prove that this map is finite, it is enough by [50, Lemma 8.7] to show that no 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is contracted to a point, that is, that $(\mathcal{Y},\mathcal{D}+\epsilon\,\mathcal{C})$ is a non-constant family of KSBA stable pairs in restriction to every 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Let $\rho_{0}$ be a codimension one cone of the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, corresponding to a 1-dimensional toric stratum $S_{\rho_{0}}$ of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. As the fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, there are two maximal cones $\rho_{1}$ and $\rho_{2}$ containing $\rho_{0}$, corresponding to the two 0-dimensional toric strata $S_{\rho_{1}}$ and $S_{\rho_{2}}$ contained in $S_{\rho_{0}}$. We will show that at least one of the fibers $(Y_{\rho_{1}},D_{\rho_{1}})$ or $(Y_{\rho_{2}},D_{\rho_{2}})$ of $(\mathcal{Y},\mathcal{D})$ over the points $S_{\rho_{1}}$ and $S_{\rho_{2}}$ is not isomorphic to the general fiber $(Y_{\rho_{0}},D_{\rho_{0}})$ of $(\mathcal{Y},\mathcal{D})$ over $S_{\rho_{0}}$.
Let $\tilde{\rho}_{0}$ be a codimension one cone of the Mori fan containing $\rho_{0}$, and $\tilde{\rho}_{1}$, $\tilde{\rho}_{2}$ be the two maximal dimensional cones of the Mori fan containing $\tilde{\rho}_{0}$ and contained in $\rho_{1}$ and $\rho_{2}$ respectively. For $i=0,1,2$, we have $\tilde{\rho}_{i}=\mathrm{Nef}(\mathcal{Z}_{i}/\overline{\mathcal{X}}_{\mathscr{P}})+\sum_{v\in J_{i}}\mathbb{R}_{\geq 0}[X^{v}]$, where $\mathcal{Z}_{i}\rightarrow\overline{\mathcal{X}}_{\mathscr{P}}$ is a partial crepant resolution, and $X^{v}$ with $v\in J_{i}$ are divisors which are contracted by a crepant resolution of $\mathcal{Z}_{i}$. Moreover, since $\tilde{\rho}_{0}$ is a codimension one face of $\tilde{\rho}_{i}$, we have a birational contraction $\mathcal{Z}_{i}\rightarrow\mathcal{Z}_{0}$, and either this map is of relative Picard rank one and $J_{i}=J_{0}$, or this map is an isomorphism and $J_{i}=J_{0}\sqcup\{V\}$ for some $V\in J_{i}\setminus J_{0}$. The later case cannot happen for both $i=1$ and $i=2$: indeed, if $J_{1}=J_{0}\sqcup\{V\}$, then $[X^{V}]\in\tilde{\rho}_{1}$ and $[X^{V}]\notin\tilde{\rho}_{0}$, and so $[X^{V}]\notin\tilde{\rho}_{2}$.
Hence, without loss of generality, we can assume that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one and $J_{1}=J_{0}$. Let $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$ be a projective crepant resolution. Denote by $H_{1}$ the face of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$ corresponding to $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$, and by $H_{0}$ the codimension one face of $H_{1}$ corresponding to $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$. By Theorem 7.12, the intersection complex of $(Y_{\rho_{1}},D_{\rho_{1}})$ (resp. $(Y_{\rho_{0}},D_{\rho_{0}})$) is given by the polyhedral decomposition $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) associated to $H_{1}$ (resp . $H_{0}$) as in §7.3.
Assume first that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is an isomorphism in codimension one. If $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ contracts a curve in the double locus of $\mathcal{Z}_{1}$, then the corresponding edge in $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$. Hence, $Y_{\rho_{1}}$, and $Y_{\rho_{0}}$ are not isomorphic, and the result is proved in this case. On the other hand, if $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ does not contract any curve in the double locus of $\mathcal{Z}_{1}$, then the exceptional locus of $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a rigid internal curve in an irreducible component of $\mathcal{Z}_{1}$, whose strict transform in $\mathcal{X}^{\prime}$ is necessarily an internal exceptional curve. It follows that $NE(\mathcal{Z}_{0}/\overline{\mathcal{X}}_{\mathscr{P}})$ is contained in the codimension one face of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$ defined by the M1 flop defined by contracting this internal exceptional curve, which is by Lemma 6.19 in contradiction with the assumption that $\rho_{0}$ is a codimension cone of the secondary fan, and so this last case actually does not occur.
Thus, it remains to treat the case where $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a divisorial contraction. Since $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one, its exceptional locus consists of a single divisor $Z_{1}^{v}$ corresponding to a vertex $v$ of $\mathscr{P}_{H_{1}}$. If an irreducible component of the boundary of $Z_{1}^{v}$ is contracted into $\mathcal{Z}_{0}$, then the corresponding edge of $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$, and so we also deduce in this case that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic. If no boundary component of $Z_{1}^{v}$ is contracted in $\mathcal{Z}_{0}$, then, as in [50, §8.2], $Z_{1}^{v}$ is necessarily a generically $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{1}$, contracted to its base $\mathbb{P}^{1}$ in $\mathcal{Z}_{0}$, and its boundary $\partial Z_{v}^{1}$ is either an irreducible nodal curve mapping 2:1 to the base, or the union of two sections intersecting in two points. If $\partial Z_{v}^{1}$ is the union of two irreducible components $D_{1}$ and $D_{2}$, consider the corresponding edges $E_{1}$ and $E_{2}$ adjacent to $v$ in $\mathscr{P}_{H_{1}}$, and choose a 2-dimensional face $\sigma_{1}$ of $\mathscr{P}_{H_{1}}$ such that $E_{1}\cup E_{2}$ are edges of the boundary of $\sigma_{1}$. Denote by $\sigma_{0}$ the same face viewed as a 2-dimensional face of $\mathscr{P}_{H_{0}}$. By Theorem 7.12, there is an irreducible component $(Y_{\sigma_{1}},\partial Y_{\sigma_{1}})$ (resp. $(Y_{\sigma_{0}},\partial Y_{\sigma_{0}})$) of $Y_{\rho_{1}}$ (resp. $Y_{\rho_{0}}$) isomorphic to the general fiber of the mirror family to the open Kulikov degeneration $\mathcal{X}_{\sigma_{1}}$ (resp. $\mathcal{X}_{\sigma_{0}}$) with tropicalization $(C(\sigma_{1}),C(\mathscr{P}_{\sigma_{1}}))$ (resp. $(C(\sigma_{1}),C(\mathscr{P}_{\sigma_{1}}))$). The vertex $v$ can be viewed as a boundary vertex of both $\mathscr{P}_{\sigma_{1}}$ and $\mathscr{P}_{\sigma_{0}}$. Since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are not contracted in $\mathcal{Z}_{1}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{1},0}$ of $\mathcal{X}_{\sigma_{1}}$ is a 00-surface. Hence, by Lemma 6.24, it corresponds to a singular point of $\partial Y_{\sigma_{1}}$. On the other hand, since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are contracted in $\mathcal{Z}_{0}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{0},0}$ of $\mathcal{X}_{\sigma_{0}}$ is a $1$-surface. Thus, by Lemma 6.24, it corresponds to a smooth point of $\partial Y_{\sigma_{0}}$. Since a singular point of $\partial Y_{\sigma_{1}}$ becomes a smooth point in $\partial Y_{\sigma_{0}}$, it follows that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic.
Finally, the case where $\partial Z^{1}_{v}$ is irreducible follows from the previous case as in [50, §8.2]. Indeed, denoting by $\sigma_{1}$ (resp. $\sigma_{0}$) the 2-dimensional face of $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) containing $v$, and cutting $\sigma_{1}$ (resp. $\sigma_{0}$) along the edges of $\mathscr{P}_{\sigma_{1}}$ (resp. $\mathscr{P}_{\sigma_{0}}$) connecting $v$ to the boundary of $\sigma_{1}$ (resp. $\sigma_{0}$), we obtain an integral affine manifold with singularities $(\tilde{\sigma}_{1},\mathscr{P}_{\tilde{\sigma}_{1}})$ (resp. $(\tilde{\sigma}_{0},\mathscr{P}_{\tilde{\sigma}_{0}})$) with the same formal properties as $(\sigma_{1},\mathscr{P}_{\sigma_{1}})$ (resp. $(\sigma_{0},\mathscr{P}_{\sigma_{0}})$) in the case where $\partial Z^{1}_{v}$ is reducible. There is a natural open Kulikov degeneration $\mathcal{X}_{\widetilde{\sigma}_{1}}$ (resp. $\mathcal{X}_{\widetilde{\sigma}_{0}}$) with tropicalization $(C(\widetilde{\sigma}_{1}),C(\mathscr{P}_{\widetilde{\sigma}_{1}}))$ (resp. $(C(\widetilde{\sigma}_{0}),C(\mathscr{P}_{\widetilde{\sigma}_{0}}))$), and the irreducible component $Y_{\tilde{\sigma}_{1}}$ (resp. $Y_{\tilde{\sigma}_{0}}$) of the general fiber of the corresponding mirror family is the normalization of $Y_{\sigma_{1}}$ (resp. $Y_{\sigma_{0}}$). As above, a singular point of $\partial Y_{\tilde{\sigma}_{1}}$ becomes smooth in $\partial Y_{\tilde{\sigma}_{0}}$. It follows that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are also non-isomorphic in this case, and this concludes the proof that the map $\mathcal{S}_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$ is finite.
To complete the proof of Theorem 8.11, it remains to show that the map $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$ is surjective. The moduli stack of log Calabi–Yau surfaces deformation equivalent to $(Y,D)$ is irreducible of dimension $Q-2$, where $Q$ is the charge of $(Y,D)$ as in [34, Definition 1.1]. Indeed, fixing a toric model $(\overline{Y},\overline{D})$, this space is covered by the torus $(\mathbb{C}^{\star})^{Q}$ of positions of the points to blow-ups on $\overline{D}$, quotiented by the action of $(\mathbb{C}^{\star})^{2}$ by toric automorphisms. Moreover, given such a log Calabi–Yau pair $(Y^{\prime},D^{\prime})$, the space of curves $C^{\prime}\in|L^{\prime}|$ such that $(Y^{\prime},D^{\prime}+\epsilon\,C^{\prime})$ is a KSBA stable pair, is open, and so dense in $|L^{\prime}|$ if non-empty. Hence, the space of KSBA stable pairs deformation equivalent to $(Y,D,L)$ is irreducible, and so the closure $\mathcal{M}_{(Y,D,L)}$ of this space in the moduli space of all KSBA stable pairs is also irreducible. On the other hand, the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, and so the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is proper. Therefore, it is enough to prove that the map $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$ is dominant.
Fix a large enough positive integer $k$ and a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $kP$ as in Lemma 8.3. By Theorem 8.10, the general fiber $(Y_{t},D_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is deformation equivalent to $(Y,D)$, and so admits a toric model $(Y_{t},D_{t})\leftarrow(\widetilde{Y}_{t},\widetilde{D}_{t})\rightarrow(\overline{Y},\overline{D})$, where $(\overline{Y},\overline{D})$ is the toric variety with momentum polytope $\overline{P}$, and the exceptional curves $E_{ij}$ of $\widetilde{Y}_{t}\rightarrow\overline{Y}$ are in one-to-one correspondence with the cut triangles $\Delta_{ij}$. Moreover, it follows from the proof of Theorem 8.10 that $(\widetilde{Y}_{t},\widetilde{D}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and that the classes of the exceptional curves $E_{ij}$ are the classes in $\widetilde{Y}_{t}$ obtained by lifting as in [71, §6.2], [83] the tropical 1-cycles defined by the cuts $\delta_{ij}$ ending at the focus-focus singularities $\widetilde{p}_{ij}$ as in §8.2. In particular, the period calculation proving [71, Proposition 6.15], based on the main result of [83], directly generalizes to our set-up and shows that the marked period of $(Y_{t},D_{t})$ marked by $\mathcal{O}_{Y_{t}}(E_{ij})$, as defined in [42], is equal to $t^{E_{ij}}$. This shows that every log Calabi–Yau surface deformation equivalent to $(Y,D,L)$ appears as a fiber $(Y_{t},D_{t},L_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Finally, given such a fiber $(Y_{t},D_{t},L_{t})$, it follows from the torus action given by Lemma 6.27 that for every $\mathbf{a}:=(a_{p})_{p\in P_{\mathbb{Z}}}\in(\mathbb{C}^{\star})^{P_{\mathbb{Z}}}$, denoting by $C_{\mathbf{a}}$ the curve in $Y_{t}$ defined by the equation $\sum_{p\in P_{\mathbb{Z}}}a_{p}\vartheta_{p}=0$, the pair $(Y_{t},D_{t}+\epsilon\,C_{\mathbf{a}})$ appears as a fiber of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Since $(\vartheta_{p})_{p\in P_{\mathbb{Z}}}$ is a basis of $H^{0}(Y_{t},L_{t})$, the set of curves of the form $C_{\mathbf{a}}$ is open dense in the linear system $|L_{t}|$. Hence, we obtain that an open dense subset of $\mathcal{M}_{(Y,D,L)}$ is contained in the image of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$, and this concludes the proof of Theorem 8.11. ∎
@@
The complete toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ with fan the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is projective.
@@
By Theorem 8.11, $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is a finite cover of the coarse KSBA moduli space $M_{(Y,D,L)}$, which is projective by [36, 69]. ∎

### v3

@@ 8.5. The Hacking–Keel–Yu conjecture
Putting together our results in previous sections, we finally prove in Theorem 8.13 below the two-dimensional case of the Hacking–Keel–Yu conjecture [53, Conjecture 1.2].
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and $(P,\mathscr{P})$ a Symington polytope with polyhedral decomposition defining a maximal degeneration of $(Y,D,L)$ as in §2.4. Let $\mathcal{M}_{(Y,D,L)}$ be the closure in the moduli space of stable pairs of the locus of stable pairs deformation equivalent to $(Y,D+\epsilon\,C)$, where $C\in|L|$ and $0<\epsilon<\!\!<1$. It is a proper Deligne–Mumford stack, with a projective coarse moduli space $M_{(Y,D,L)}$ by [38, 75].
Lemma 8.12.
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and $(P,\mathscr{P})$ a Symington polytope with polyhedral decomposition defining a maximal degeneration of $(Y,D,L)$ as in §2.4. Then, $M_{(Y,D,L)}$ is an irreducible projective variety of dimension
$\dim M_{(Y,D,L)}=Q+|P_{\mathbb{Z}}|-3\,$,
where $Q$ is the charge of $(Y,D)$ as in [36, Definition 1.1], that is, the number of interior blow-ups in a toric model of $(Y,D)$, and $|P_{\mathbb{Z}}|$ is the number of integral points in $P$.
Proof.
By the Torelli theorem for log Calabi–Yau surfaces [45, Theorem 6.1], the moduli stack of log Calabi–Yau surfaces deformation equivalent to $(Y,D)$ is irreducible of dimension Q2Q-2. Moreover, given such a log Calabi–Yau pair $(Y^{\prime},D^{\prime})$, the space of curves $C^{\prime}\in|L^{\prime}|$ such that $(Y^{\prime},D^{\prime}+\epsilon\,C^{\prime})$ is a KSBA stable pair, is open, and so dense in $|L^{\prime}|$ if non-empty. By the dimension count following [31, Conjecture 6.3.7], we obtain
$\dim H^{0}(Y^{\prime},L^{\prime})=|P_{\mathbb{Z}}|\,,\quad\text{so}\quad\dim|L^{\prime}|=|P_{\mathbb{Z}}|-1\,.$
Thus, the space of KSBA stable pairs deformation equivalent to $(Y,D,L)$ is irreducible of dimension $Q+|P_{\mathbb{Z}}|-3$. Consequently, its closure $\mathcal{M}_{(Y,D,L)}$ in the moduli space of all KSBA stable pairs is also irreducible of dimension $Q+|P_{\mathbb{Z}}|-3$. ∎
@@
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and $(P,\mathscr{P})$ a Symington polytope with polyhedral decomposition defining a maximal degeneration of $(Y,D,L)$ as in §2.4. Let $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\to\Delta$ be the corresponding semistable mirror as in §4.3. Then, there exists a finite surjective morphism $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow M_{(Y,D,L)}$, from the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ with fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$, to the coarse KSBA moduli space $M_{(Y,D,L)}$.
@@
By Theorem 7.7, the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is the base of a family of KSBA stable log Calabi–Yau surfaces which by Theorem 8.11 are generically deformation equivalent to $(Y,D,L)$. Hence, by the universal property of the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$, there exists a corresponding morphism of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$, and so an induced morphism $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to M_{(Y,D,L)}$ at the level of the coarse moduli spaces. By Theorem 7.8, the KSBA stable family over $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is non-constant in restriction to every 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$, hence, no 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is contracted to a point in $M_{(Y,D,L)}$. By [53, Lemma 8.7], this implies that $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to M_{(Y,D,L)}$ is a finite morphism. By Lemma 3.11 and Lemma 8.12, $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ and $M_{(Y,D,L)}$ are both proper irreducible varieties, and $\dim\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\geq Q+|P_{\mathbb{Z}}|-3=\dim M_{(Y,D,L)}$. Hence, the finite morphism $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to M_{(Y,D,L)}$ is necessarily surjective, and this completes the proof of Theorem 8.13. ∎
@@
The complete toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ with fan the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$ is projective of dimension $Q+|P_{\mathbb{Z}}|-3$.
@@
By Theorem 8.13, $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is a finite cover of the coarse KSBA moduli space $M_{(Y,D,L)}$, which is projective by [38, 75], and so $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is also projective. Moreover, $\dim S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}=\dim M_{(Y,D,L)}=Q+|P_{\mathbb{Z}}|-3$ by Lemma 8.12. ∎